OpenStudy (anonymous):

Geometry help anyone? Thank you so much. :) What is the lateral area of a triangular prism with base edges all equaling 2 millimeters and a height of 7 millimeters? A. 7 mm² B. 14 mm² C. 28 mm² D. 42 mm² I just don't seem to understand this lesson, I'm usually pretty okay at Geometry, this lesson has just been throwing me off guard. I assume I would use the Lateral Area formula?

OpenStudy (anonymous):

If it helps anyone, the formula I have for Lateral Area is LA= PH and that is specifically for prisms.

OpenStudy (anonymous):

The P being perimeter and the H being height.

OpenStudy (bahrom7893):

how many edges does a triangular prism have?

OpenStudy (anonymous):

well, a rectangular prism has 2 bases, so it would have 3 base edges then, correct?

OpenStudy (bahrom7893):

3 base edges on each side, that makes 6 base edges in total and plus 3 edges on the sides. So 9 right? 

OpenStudy (anonymous):

@bahrom7893 I do believe that is correct, although I have someone else telling me that the answer is C. 28. Although, at this point, i'm just confused- so if you could tell me what you think of that, i'd be grateful.

OpenStudy (bahrom7893):

The surface area of a prism = 2 × area of base + perimeter of base × H

OpenStudy (bahrom7893):

Area of base: p=(2+2+2)/2=3 A=sqrt(p(p-a)(p-b)(p-c))=sqrt(3*1*1*1)=sqrt(3)

OpenStudy (bahrom7893):

hmm I'm lost. @amistre64 can you take a look at this. That sqrt(3) is bothering me

OpenStudy (amistre64):

lateral area is just the area of the sides without the bases

OpenStudy (bahrom7893):

OHHHH NO WONDER!!!

OpenStudy (amistre64):

3(2*7)